To solve the break even EBIT

4. Breakeven even EBIT James Corporationis comparing two different capital structures: an all equity plan (plan I) and levered plan (plan II). Under plan I , thecompany would have 160,000
shares of stuck outstanding. Under plan (plan II), there is would be 80,000 shares of stock outstanding and $2.8 million in debt outstanding. The interest rate on thedebt is 8%,
and there are no taxes.

a. If EBIT is $ 350,000, which plan will result in the higher EPS?
b. If EBIT is $ 500,000, which plan will result in the higher EPS?
c. What is the break-even EBIT?
Answer #1
a. EBIT = $350,000, Equity Shares = 160,000, Debt = 0 so Int = 0
So EPS = (EBIT - Int)/No of Equity shares = $350,000/160,000 = $2.1875

b. No of shares = 80,000
Debt = $2.8M. Int Rate = 8%
SO Int on Debt = 8%*$2.8M = $224,000
So EPS = (EBIT - Int)/No of shares = ($500,000 - $224,000)/80,000 = $276,000/80,000=$3.45

c. At Break even EBIT, EPS = 0. This is the EPS wherein there is no profit or loss.
As nt = $224,000, So EBIT = $224,000
This will give EPS = 0 as EBIT - Int = 0
SO Brk Even EBIT = $224,000